OpenStudy (anonymous):

I'm having some trouble with this question. will post my thought process in a bit! For the thermal decomposition of hydrogen sulfide, H2S(g) ↔ 2H2(g) + S2(g), K = 2.2x10-4 at 1,400 K. A sample of gas in which [H2S] = 6.00 M is heated to the decomposition point in a sealed vessel. After equilibrium is reached, what is [H2S]? You can assume that no H2 or S2 was present in the original sample. A. 3.22 M B. 5.75 M C. 4.45 M D. 1.77 M E. 5.10 M

OpenStudy (anonymous):

First I balanced the equation: 2H2S(g) ↔ 2H2(g) + S2(g) and use the ICE table, which I would get 2.2*10^-4=(x)*(2x^2)/(6-2x)^2 solving for x, I would get 0.153, plugging it back into the equation is 6-2(0.153)=5.69M but no choices are that concentration. the closest is B 5.75, but I think this question is harder than I think it is.

OpenStudy (anonymous):

\[K = 2.2E-4 = \frac{ [H2(g)]^2[S2(g)] }{ [H2S(g)] } = \frac{ [2x]^2[x] }{ [6-2x]^2 } = \frac{ [4x^2][x] }{ [6-2x]^2 } = \frac{ [4x^3] }{ [6-2x]^2 }\] is what i have so far

OpenStudy (anonymous):

gosh this equation thingy takes forever to do

OpenStudy (anonymous):

I also tried using the free energy side of it. deltaG=-RTln(K) free energy is always 0 at equilibrium so 0=-8.314*1400*ln((x)*(2x^2)/(6-2x)^2) which solving for x would be 1.587, plugging that into the ice table; it would be 6-2(1.587)=2.83M which again is none of the choices

OpenStudy (anonymous):

Thanks for staying with me, you can use just the easier way to enter equations if you want to :)

OpenStudy (anonymous):

wait

OpenStudy (anonymous):

how did you solve for x? I've having a hard time doing that lol

OpenStudy (anonymous):

i tried square rooting both sides and got 0.014 = \[\sqrt{x}\]/6-2x

OpenStudy (anonymous):

ops i mean, \[0.014 =\frac{ 2x \sqrt{x} }{ 6-2x }\]

OpenStudy (anonymous):

ya i was going to say that the equation in the problem wasn't balanced lol

OpenStudy (anonymous):

But @premedhelp , x can't be 0.153 b/c if you plug back 0.153 into the equilibrium expression, you don't get the same equilibrium constant of 2.2E-4

OpenStudy (anonymous):

thats what I was thinking, but doesnt the constant change when you heat it up to the decomposition point? i think thats what im confused about

OpenStudy (anonymous):

and neither can it be 1.587 b/c it doesn't give you 2.2E-4

OpenStudy (anonymous):

Oh snaps, yes you're definitely right. The eq constant's value changes with temperature, but we're not told the final temperature so we can't find the new eq constant

OpenStudy (anonymous):

my other question is if 6M is even the initial concentration of H2S though

OpenStudy (anonymous):

yep, has been bugging me an hour

OpenStudy (anonymous):

ya it has to be. it can't be the final eq concentration b/c that would be the answer then lol

OpenStudy (anonymous):

Look, i think we have our equilibrium expressions set up correctly. I'm thinking that where we're going wrong is actually solving for x

OpenStudy (anonymous):

that's all we have to do, solve for x properly

OpenStudy (anonymous):

http://www.wolframalpha.com/input/?i=2.2*10^-4%3D%28x%29*%282x^2%29%2F%286-2x%29^2 here could be our reference, just cant get step by step

OpenStudy (anonymous):

no you typed it up wrong

OpenStudy (anonymous):

it should be (x)(2x)^2 or (x)(4x^2) or just 4x^3

OpenStudy (anonymous):

YAY after we type it in correctly, we'll solve this problem correctly YES!

OpenStudy (anonymous):

and just to check, we can plug x = 0.122 into the equilibrium expression and we DO get 2.2E-4 :)

OpenStudy (anonymous):

Now we can solve for the equilibrium concentration of H2S(g), which is represented by 6-2x

OpenStudy (anonymous):

which is 5.75M!

OpenStudy (anonymous):

so we can substitute our new x of 0.122 into 6-2x and we get 5.756

OpenStudy (anonymous):

but i would still like to know how we can solve for x manually

OpenStudy (anonymous):

YAY! WE FRIGGING DID IT LOL

OpenStudy (anonymous):

high five :)

OpenStudy (anonymous):

;)

OpenStudy (anonymous):

ya me too :(

OpenStudy (anonymous):

stay on for a bit, and ill try to figure it out :)

OpenStudy (anonymous):

but damn...that wolf ram alpha tho... lol i've never used that before. That program is cool

OpenStudy (anonymous):

haha message me it, i'm going to eat right now :) Good luck

OpenStudy (anonymous):

i know right! alrighty, eat up! :)

OpenStudy (anonymous):

I need some...."food for thought" haha

OpenStudy (anonymous):

haha, you're right! If you want to make a friend, I'll add you on fb?

OpenStudy (anonymous):

ya sure

OpenStudy (anonymous):

i've sent you my fb link to ur message inbox